OpenStudy (anonymous):

Solve the following system. Show all your work. 3x-2y+2z=30 -x+3y-4z=-33 2x-4y+3z=42 Please help me I really need your help can you tell me how to do it too I dont get it... I need it ASAP

OpenStudy (anonymous):

x

OpenStudy (anonymous):

Thank you soo much

OpenStudy (mathstudent55):

You can use elimination to eliminate one variable. Then you will have a system of 2 equations with 2 unknowns to solve.

OpenStudy (mathstudent55):

Look at equations 1 and 2. Can you eliminate x?

OpenStudy (anonymous):

yes 2x

OpenStudy (anonymous):

Okay let me do that real quick

OpenStudy (mathstudent55):

No. 2x is not eliminating x. Equation 1 has 3x. Equation 2 has -x. Ok so far?

OpenStudy (anonymous):

2x+y-2z-3

OpenStudy (anonymous):

Is what you get when you eliminate am I correct?

OpenStudy (mathstudent55):

If you multiply equation 2 by 3, you will have -3x. Then when you add the new equation 2 to equation 1, x will be eliminated.

OpenStudy (anonymous):

Oh multiply K

OpenStudy (mathstudent55):

Each time you add equations, you need to eliminate a variable.

OpenStudy (anonymous):

I really dont understand this..

OpenStudy (anonymous):

And I pay attention in class 2...

OpenStudy (anonymous):

It is just soo confusing

OpenStudy (mathstudent55):

We'll do it step by step. Here are the two equations 1 and 2 in their original form. Eq.1 3x-2y+2z=30 Eq. 2 -x+3y-4z=-33

OpenStudy (mathstudent55):

Now we keep equation 1 as is and we multioply the entire equation 2 by 3, and write it below eq. 1.

OpenStudy (anonymous):

Oh by 3 Okay

OpenStudy (mathstudent55):

Eq.1 3x - 2y + 2z = 30 New Eq. 2 -3x + 9y - 12z= -99

OpenStudy (mathstudent55):

Now we add them together. Notice, the x terms cancel out.

OpenStudy (anonymous):

Yes so the equation afterwards will be 7y -10z=-69

OpenStudy (mathstudent55):

Eq.1 3x - 2y + 2z = 30 New Eq. 2 -3x + 9y - 12z= -99 -------------------------------- add 7y - 10z = -69

OpenStudy (mathstudent55):

Correct.

OpenStudy (mathstudent55):

Now we need to pick a different pair of equations and again eliminate x.

OpenStudy (mathstudent55):

Let's use equations 2 and 3.

OpenStudy (anonymous):

Okay

OpenStudy (mathstudent55):

Eq. 2 -x+3y-4z=-33 Eq. 3 2x-4y+3z=42 By multiplying Eq. 2 by 2, the x's will be eliminated after adding the two eqs togeher.

OpenStudy (anonymous):

so it will be -y-z=9

OpenStudy (mathstudent55):

You must multiply the entire eq 2 by 2. New Eq. 2 -2x + 6y - 8z = -66 Eq. 3 2x - 4y + 3z = 42 ------------------------------ add 2y - 5z = -24

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

And than after that you get both of the answers that you simplified and and them together

OpenStudy (anonymous):

So 7y-10z=-69 2y-5z =-24

OpenStudy (mathstudent55):

Now take the two equations with x eliminated, and you have a system of 2 equations in 2 unknowns.

OpenStudy (mathstudent55):

Correct. Since we are already using elimination, let's continue with that method. 7y - 10z = -69 2y - 5z = -24 Look at -10z and -5z. If we multiply the seciond eq by -2, the z term will become 10z which when added to -10z of the eq above will eliminate z.

OpenStudy (anonymous):

This is the part I dont get...

OpenStudy (mathstudent55):

We are using elimination to solve a system of 2 equations n 2 unknowns. Keep eq. as is 7y - 10z = -69 Multiplied by -2 -4y + 10z = 48 ------------------- add 3y = -21 y = -7

OpenStudy (anonymous):

So you multiplied the bottom equation by -2 correct?

OpenStudy (mathstudent55):

yes

OpenStudy (anonymous):

And what do I do after that I found y but after that do I substitute it in with the other equations?

OpenStudy (mathstudent55):

Now that we have y, we use one of the two equation that have only y and z to find z. Let's use the first one: 7y - 10z = -69, using y = -7: 7(-7) - 10z = -69 Solve for z

OpenStudy (anonymous):

-49-10z=-69

OpenStudy (anonymous):

Than you add -49 into -69 but the -49 becomes a positive

OpenStudy (mathstudent55):

7(-7) - 10z = -69 -49 - 10z = -69 -10z = -20 z = 2

OpenStudy (anonymous):

Thanks!

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

And you get x by doing what...

OpenStudy (mathstudent55):

Now that we have y and z, we go back to one of the original 3 equations that has x, y, and z. We plug in the known values of y and z, and we solve for x.

OpenStudy (mathstudent55):

Let's use the first original equation: 3x - 2y + 2z = 30

OpenStudy (mathstudent55):

3x - 2(-7) + 2(2) = 30 Now you have one equation with only one variable, so you can solve for x.

OpenStudy (anonymous):

3x+18=30 x=4

OpenStudy (anonymous):

x=4

OpenStudy (mathstudent55):

3x + 14 + 4 = 30 3x + 18 = 30 3x = 12 x = 4

OpenStudy (anonymous):

Because 3x=12 divide by 3

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

:D So y=7, z=, 2, and x=4

OpenStudy (anonymous):

Thank you soooo much!!!

OpenStudy (mathstudent55):

The solution is: x = 4; y = -7; z = 2

OpenStudy (mathstudent55):

You are very welcome.